The planet Mars has a mass of 6.1 × 1023 kg
and radius of 3.1 × 106 m.
What is the acceleration of an object
in free fall near the surface of Mars?
The value of the gravitational constant is
6.67259 × 10−11 N · m2/kg2.
Answer in units of m/s2.
1
Expert's answer
2011-01-18T15:44:57-0500
The gravitation acceleration can be written as
g = GM/R2 = 6.67259 × 10−11 * 6.1 × 1023 / (3.1 × 106)2 = 4.24 m/s2.